### 27 Mar 2026 v2

# Unimodular rows over affine algebras over algebraic closure of a finite field

Abstract.
In this article, we prove that if $R$ is an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$ and $1/(d-1)!\in R,$ then any unimodular row over $R$ of length $d$ can be mapped to a factorial row by elementary transformations.
020 Mathematics Subject Classification:13C10
Unimodular row, nice group structure
1. Introduction
All the rings are assumed to be commutative noetherian with unity $1$.
One can say that the study of unimodular rows came in the picture with the Serre’s question about projective modules. In [9], Serre asked whether finitely generated projective modules over $k[X_{1},\ldots,X_{n}]$ free, where $k$ is a field. This question of Serre was answered by Quillen [6] and Suslin [11] independentaly. This is now known as Quillen-Suslin theorem. Due to the work of Grothendieck, one knows that every finitely generated projective module over $k[X_{1},\ldots,X_{n}]$ is stably free. To a stably free module one can associate a unimodular row and vice versa. And the connection between these two objects is that a stably free module is free if and only if the corresponding unimodular row is completable to an invertible matrix. Therefore question about freeness of a projective module boils down to the question about completion of certain unimodular row. This shows that study of unimodular rows is quite important in classical $K$-theory.
In [15], R.G. Swan and J. Towber showed that if $(a^{2},b,c)\in Um_{3}(R)$ then it can be completed to an invertible matrix over $R.$ This result of Swan and Towber was generalised by Suslin in [13] who showed that if $(a_{0}^{r!},a_{1},\ldots,a_{r})\in Um_{r+1}(R)$ then it can be completed to an invertible matrix. Rao also studied the unimodular rows over $R[X]$ where $R$ is a local ring in [7], [8]. In [10], author studied the unimodular rows over finitely generated rings over $\mathbb{Z}$ and proved that if $R$ is a finitely generated ring over $\mathbb{Z}$ of dimension $d,d\geq 2,\frac{1}{d!}\in R$, then any unimodular row over $R[X]$ of length $d+1$ can be mapped to a factorial row by elementary transformations.
In [3, Theorem 7.5], Fasel, Rao and Swan proved that if $R$ is a normal affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$ and $1/(d-1)!\in R,$ then any unimodular row over $R$ of length $d$ can be mapped to a factorial row by elementary transformations. In this article we remove their normality condition. In particular, we prove the following :
Theorem 1.1.
Let $R$ be an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$ and $1/(d-1)!\in R.$ Let $v=(v_{1},\ldots,v_{d})\in\mbox{\rm Um}_{d}(R).$ Then there exists $\varepsilon\in\mbox{\rm E}_{d}(R)$ such that
$(v_{1},\ldots,v_{d})\varepsilon=(w_{1},w_{2},\ldots,w_{d}^{(d-1)!})$
for some $(w_{1},\ldots,w_{d})\in\mbox{\rm Um}_{d}(R).$ In particular, $\mbox{\rm Um}_{d}(R)=e_{1}\mbox{\rm SL}_{d}(R).$
In this article, we consider the action of symplectic matrices on unimodular rows and prove the following :
Theorem 1.2.
Let $R$ be an affine algebra of dimension $d$ over $\overline{\mathbb{F}}_{p}$ and $d\equiv 2~(\mbox{mod}~4).$ Let $v\in\mbox{\rm Um}_{d}(R).$ Assume that $\mbox{\rm SL}_{d+1}(R)\cap\mbox{\rm E}(R)=\mbox{\rm E}_{d+1}(R).$ Then $v$ is completable to a symplectic matrix.
We also studied the problem of injective stability for affine algebras over $\overline{\mathbb{F}}_{p}$ and proved the following :
Theorem 1.3.
Let $R$ be an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$. Let $\alpha\in\mbox{\rm SL}_{d}(R)\cap\mbox{\rm E}_{d+1}(R).$ Then $\alpha$ is isotopic to the identity matrix.
2. Preliminaries
In this section, we give some basic definitions and collect some known results which will be used later in the article.
2.1. Basic definitions and notations
Definition 2.1.
A row $v=(a_{0},\ldots,a_{r})\in R^{r+1}$ is said to be unimodular if there is a $w=(b_{0},\ldots,b_{r})\in R^{r+1}$ with $\langle v,w\rangle=\Sigma_{i=0}^{r}a_{i}b_{i}=1$ and $Um_{r+1}(R)$ will denote the set of unimodular rows (over $R$) of length $r+1$.
The elementary linear group $E_{r+1}(R)$ acts on the rows of length $r+1$ by right multiplication. Moreover, this action takes unimodular rows to unimodular rows : $\frac{Um_{r+1}(R)}{E_{r+1}(R)}$ will denote the set of orbits of this action; and we shall denote by $[v]$ the equivalence class of a row $v$ under this equivalence relation.
Definition 2.2.
A unimodular row $v\in Um_{n}(R)$ is said to be completable if there exists a matrix $\alpha\in GL_{n}(R)$ such that $v=e_{1}\alpha.$
Notation 2.3.
Let $\psi_{1}=\begin{bmatrix}0&1\\ -1&0\\ \end{bmatrix},~~\psi_{n}=\psi_{n-1}\perp\psi_{1}$ for $n>1.$
Notation 2.4.
Let $\sigma$ be the permutation of the natural numbers given by $\sigma(2i)=2i-1$ and $\sigma(2i-1)=2i$.
Definition 2.5.
Symplectic group $Sp_{2m}(R):$ the group of all $2m\times 2m$ matrices $\{\alpha\in GL_{2m}(R)~\mid\alpha^{t}\psi_{m}\alpha=\psi_{m}\}$.
Definition 2.6.
Elementary symplectic group $ESp_{2m}(R)$: We define for $1\leq i\neq j\leq 2m,~z\in R,$
$se_{ij}(z)=\begin{cases}I_{2m}+zE_{ij},~~~\textit{if}~i=\sigma(j);\\ I_{2m}+zE_{ij}-(-1)^{i+j}zE_{\sigma(j)\sigma(i)},~~~\textit{if}~i\neq\sigma(j).\par\end{cases}$
It is easy to verify that all these matrices belong to $Sp_{2m}(R)$. We call them the elementary symplectic matrices over $R$. The subgroup generated by them is called the elementary symplectic group and is denoted by $ESp_{2m}(R)$.
Definition 2.7.
Let $I$ be an ideal of a ring $R.$ A unimodular row $v\in Um_{n}(R)$ which is congruent to $e_{1}$ modulo $I$ is called unimodular relative to ideal $I.$ Set of unimodular rows relative to ideal $I$ will be denoted by $Um_{n}(R,I).$
Let $I$ be an ideal of a ring $R$. We shall denote by $GL_{n}(R,I)$ the kernel of the canonical mapping $GL_{n}(R)\longrightarrow GL_{n}(\frac{R}{I}).$ Let $SL_{n}(R,I)$ denotes the subgroup of $GL_{n}(R,I)$ of elements of determinant $1$.
Let $I$ be an ideal of a ring $R$. We shall denote by $Sp_{2m}(R,I)$ the kernel of the canonical mapping $Sp_{2m}(R)\longrightarrow Sp_{2m}(\frac{R}{I}).$
Definition 2.8.
$\bf{The~relative~groups~E_{n}(I),~E_{n}(R,I):}$} Let $I$ be an ideal of $R$. The elementary group $E_{n}(I)$ is the subgroup of $E_{n}(R)$ generated as a group by the elements $e_{ij}(x),~x\in I,~1\leq i\neq j\leq n.$
The relative elementary group $E_{n}(R,I)$ is the normal closure of $E_{n}(I)$ in $E_{n}(R)$.
Definition 2.9.
$\bf{The~relative~groups~ESp_{2m}(I),~ESp_{2m}(R,I):}$} Let $I$ be an ideal of $R$. The elementary symplectic group $ESp_{2m}(I)$ is the subgroup of $ESp_{2m}(R)$ generated as a group by the elements $se_{ij}(x),~x\in I,~1\leq i\neq j\leq 2m.$
The relative elementary symplectic group $ESp_{2m}(R,I)$ is the normal closure of $ESp_{2m}(I)$ in $ESp_{2m}(R)$.
2.2. The Suslin matrices
First recall the Suslin matrix $S_{r}(v,w)$. These were defined by Suslin in (([13, Section 5]).). We recall his inductive process : Let $v=(a_{0},v_{1}),w=(b_{0},w_{1})$, where $v_{1},w_{1}\in M_{1,r}(R)$. Set $S_{0}(v,w)=a_{0}$ and set
$S_{r}(v,w)=\begin{bmatrix}a_{0}I_{2^{r-1}}&S_{r-1}(v_{1},w_{1})\\ -S_{r-1}(w_{1},v_{1})^{T}&b_{0}I_{2^{r-1}}\\ \end{bmatrix}$.
The process is reversible and given a Suslin matrix $S_{r}(v,w)$ one can recover the associated rows $v,w$, i.e. the pair $(v,w)$.
In [12], Suslin proves that if $\langle v,w\rangle=v\cdot w^{T}=1$, then by row and column operations one can reduce $S_{r}(v,w)$ to a matrix $\beta_{r}(v,w)$ of size $r+1$ whose first row is $(a_{0},a_{1},a_{2}^{2}\cdots,a_{r}^{r})$. This in particular proves that rows of such type can be completed to an invertible matrix of determinant $1.$ We call $\beta_{r}(v,w)$ to be a compressed Suslin matrix.
2.3. Some assorted results
We state two results from [16, Corollary 17.3, Corollary 18.1, Theorem 18.2].
Theorem 2.1.
Let $R$ be an affine $C$-algebra of dimension $d\geq 2$, where $C$ is either a subfield $F$ of $\overline{\mathbb{F}}_{p}$ or $C=\mathbb{Z}$. Then
  • If $d=2,$ then ${\mbox{\rm E}}_{3}(R)$ acts transitively on $\mbox{\rm Um}_{3}(R).$
  • If $d\geq 3,$ then $\mbox{sr}(R)\leq d.$ As a consequence, ${\mbox{\rm E}}_{d+1}(R)$ acts transitively on $\mbox{\rm Um}_{d+1}(R).$
We will state the part of Swan’s Bertini needed in our proof [14, Theorems 1.3, 1.4] (see also [5, Theorem 2.3]).
Theorem 2.2.
Let $R$ be a geometrically reduced affine ring of dimension $d$ over an infinite field $k$ and $(a,a_{1},\ldots,a_{r})\in\mbox{\rm Um}_{r+1}(R)$. Then there exist $b_{1},\ldots,b_{r}\in R$ such that if $a^{\prime}=a+a_{1}b_{1}+\ldots+a_{r}b_{r}$, then $R/(a^{\prime})$ is a reduced affine algebra of dimension d1d-1 which is smooth at all smooth points of $R$.
We note a result of Fasel, Rao and Swan in [14, Theorems 7.5].
Theorem 2.3.
Let $R$ be a smooth affine algebra of dimension $d\geq 3$ over an algebraically closed field $k$ and $\mbox{gcd}(d-1)!,\mbox{char}(k))=1.$ Let $v=(v_{1},\ldots,v_{d})\in\mbox{\rm Um}_{d}(R).$ Then there exists $\varepsilon\in E_{d}(R)$ such that
$v\varepsilon=(w_{1},w_{2},\ldots,w_{d}^{(d-1)!})$
for some $(w_{1},w_{2},\ldots,w_{d})\in\mbox{\rm Um}_{d}(R).$
Next we note a result of Vaserstein [16, Lemma 5.5].
Lemma 2.4.
Let $R$ be a ring. Then for any $m\geq 1,$ ${\mbox{\rm E}}_{2m}(R)e_{1}=({\mbox{\rm Sp}}_{2m}(R)\cap{\mbox{\rm E}}_{2m}(R))e_{1}.$
Remark 2.5.
It was observed in [2, Lemma 2.13] that Vaserstein’s proof actually shows that ${\mbox{\rm E}}_{2m}(R)e_{1}={\mbox{\rm ESp}}_{2m}(R)e_{1}.$
We recall a result of Basu, Chattopadhyay and Rao [1, Lemma 2.13].
Lemma 2.6.
Let $R$ be a ring essentially of dimension $d\geq 1$ and $S_{r}(v,w)\in\mbox{\rm SL}_{2^{r}(R,I)},2^{r}\geq d+2$ for $v,w\in\mbox{\rm Um}_{r+1}(R,I)$ satisfying $\langle v,w\rangle=1.$ Let us assume that $r\equiv 1~(\mbox{mod}~4).$ Then there exists $\varepsilon_{J_{r}}\in\mbox{\rm ESp}_{J_{r}}(R,I)$ such that $S_{r}(v,w)\varepsilon_{J_{r}}=(I_{2^{r}-k}\perp\gamma)\varepsilon$ for some $\gamma\in\mbox{\rm Sp}_{k}(R,I)$ and $\varepsilon\in\mbox{\rm E}_{2^{r}}(R,I).$ Here $k=d+1$ if $d$ is odd and $k=d$ if $d$ is even.
3. Completion to a special linear matrix
In this section we prove that unimodular rows of length $d$ over $R$ can be elementarily mapped to a factorial row when R is an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$ and $1/(d-1)!\in R.$ We also prove this result in the relative case.
Theorem 3.1.
Let $R$ be an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$ and $1/(d-1)!\in R.$ Let $v=(v_{1},\ldots,v_{d})\in\mbox{\rm Um}_{d}(R).$ Then there exists $\varepsilon\in\mbox{\rm E}_{d}(R)$ such that
$(v_{1},\ldots,v_{d})\varepsilon=(w_{1},w_{2},\ldots,w_{d}^{(d-1)!})$
for some $(w_{1},\ldots,w_{d})\in\mbox{\rm Um}_{d}(R).$ In particular, $\mbox{\rm Um}_{d}(R)=e_{1}\mbox{\rm SL}_{d}(R).$
$\bf Proof:$}} In view of [7, Remark 1.4.3],we may assume that $R$ is reduced. Let $J$ be the ideal defining singular locus of $R$. Since $R$ is a reduced ring, $\mbox{ht}(J)\geq 1.$ Let $\overline{R}=R/J.$ Thus $\mbox{dim}(\overline{R})\leq d-1$ and $\overline{v}=(\overline{v}_{1},\ldots,\overline{v}_{d})\in\mbox{\rm Um}_{d}(\overline{R}).$ In view of Theorem 2.1, $\mbox{sr}(\overline{R})\leq d-1.$ Thus there exist $\overline{\lambda}_{2},\ldots,\overline{\lambda}_{d}\in\overline{R}$ such that $(\overline{v}_{2}^{\prime},\ldots,\overline{v}_{d}^{\prime})\in\mbox{\rm Um}_{d-1}(\overline{R}),$ where $\overline{v}_{i}^{\prime}=\overline{v}_{i}+\overline{\lambda_{i}v_{i}},2\leq i\leq d.$ Note that $(\overline{v}_{1},\overline{v}_{2},\ldots,\overline{v}_{d})]=[(\overline{v}_{1},\overline{v}_{2}^{\prime},\ldots,\overline{v}_{d}^{\prime})].$
Since $(\overline{v}_{2}^{\prime},\ldots,\overline{v}_{d}^{\prime})\in\mbox{\rm Um}_{d-1}(\overline{R}),$ we can elementarily transform $(\overline{v}_{1},\overline{v}_{2}^{\prime},\ldots,\overline{v}_{d}^{\prime})$ to $(1,0,\ldots,0).$ Thus we may assume that $(v_{1},\ldots,v_{d})\equiv e_{1}~\mbox{mod}~J.$ Now in view of the Theorem 2.2, we add multiples of $v_{2},\ldots,v_{d}$ to $v_{1}$ to transform it to $v_{1}^{\prime}$ such that $R/(v_{1}^{\prime})$ is smooth at the smooth points of $R.$ Since $v_{1}^{\prime}\equiv 1~\mbox{mod}~J$, $R/(v_{1}^{\prime})$ is smooth of dimension d1.d-1.
Thus in view of the Theorem 2.3, there exists $\overline{\varepsilon}\in\mbox{\rm E}_{d-1}(\overline{R})$ such that $(\overline{v}_{2},\ldots,\overline{v}_{d})\overline{\varepsilon}=(\overline{w}_{2},\ldots,\overline{w}_{d}^{(d-1)!}).$ Let $\varepsilon$ be a lift of $\overline{\varepsilon}$ and upon making appropriate elementary transformations we get
$(v_{1},v_{2},\ldots,v_{d})\varepsilon=(w_{1},w_{2},\ldots,w_{d}^{(d-1)!}).$
[…]
Lemma 3.2.
Let $R$ be an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$ and let $v\in\mbox{\rm Um}_{d}(R,I)$ for some ideal $I\subset R.$ Let $J$ denotes the ideal defining singular locus of $R$. Then there exists $\varepsilon\in\mbox{\rm E}_{d}(R,I)$ such that $v\varepsilon=(u_{1},u_{2},\ldots,u_{d})$ with $u_{1}\equiv 1~\mbox{mod}~(I\cap J).$
$\bf Proof:$}} Let $v=(v_{1},v_{2},\ldots,v_{d})\in\mbox{\rm Um}_{d}(R,I)$ and $\overline{R}=R/J.$ We have $\overline{v}=(\overline{v_{1}},\overline{v_{2}},\ldots,\overline{v_{d}})\in\mbox{\rm Um}_{d}(\overline{R},\overline{I}).$ In view of the Theorem 2.1, $\mbox{sr}(\overline{R})\leq d-1.$ Thus upon adding multiples of $\overline{v_{d}}$ to $\overline{v_{1}},\overline{v_{2}},\ldots,\overline{v_{d-1}}$, we may assume that $(\overline{v_{1}},\overline{v_{2}},\ldots,\overline{v_{d-1}})\in\mbox{\rm Um}_{d-1}(\overline{R},\overline{I}).$ Let $(\overline{w_{1}},\overline{w_{2}},\ldots,\overline{w_{d-1}})\in\mbox{\rm Um}_{d-1}(\overline{R},\overline{I})$ be such that $\sum_{i=1}^{d-1}\overline{v_{i}}\overline{w_{i}}=\overline{1}.$ Let
$\overline{\varepsilon_{1}}=\begin{bmatrix}\overline{1}&\overline{0}&\ldots&(1-\overline{v_{d}})\overline{w_{1}}\\ \overline{0}&\overline{1}&\ldots&(1-\overline{v_{d}})\overline{w_{2}}\\ \vdots&&&\vdots\\ \overline{0}&\overline{0}&\ldots&(1-\overline{v_{d}})\overline{w_{d-1}}\\ \overline{0}&\overline{0}&\ldots&\overline{1}\\ \end{bmatrix}\in\mbox{\rm E}_{d}(\overline{R}).$
[…]
Theorem 3.3.
Let $R$ be an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$ and let $v\in\mbox{\rm Um}_{d}(R,I)$ for some ideal $I\subset R.$ Then there exists $\varepsilon\in\mbox{\rm E}_{d}(R,I)$ such that
$v\varepsilon=(w_{1},\ldots,w_{d}^{(d-1)!})$
for some $(w_{1},\ldots,w_{d})\in\mbox{\rm Um}_{d}(R,I).$
$\bf Proof:$}} Let $v=(v_{1},\ldots,v_{d})\in\mbox{\rm Um}_{d}(R,I)$ and $J$ be the ideal defining singular locus of $R.$ In view of the Lemma 3.2, we may assume that $v_{1}\equiv 1~\mbox{mod}~(I\cap J).$ Let $v_{1}=1-\lambda$ for some $\lambda\in I\cap J.$ Note that $(v_{1},\ldots,v_{d})\underset{\mbox{\rm E}_{d}(R,I)}{\sim}(v_{1},\lambda v_{2},\ldots,\lambda v_{d}).$ Now in view of the Theorem 2.2, we add multiples of $\lambda v_{2},\ldots,\lambda v_{d}$ to $v_{1},$ to transform it to $v_{1}^{\prime}$ such that $R/(v_{1}^{\prime})$ is smooth outside the singular set of $R$. Since $v_{1}^{\prime}\equiv 1~\mbox{mod}~J$, $R/(v_{1}^{\prime})$ is a smooth affine algebra of dimension d1.d-1.
Let $v_{1}^{\prime}=1-\eta$ for some $\eta\in I\cap J.$ Note that in $\overline{R}=R/(v_{1}^{\prime}),$ we have $\overline{\eta}=\overline{1}.$ In view of the Theorem 2.3, there exists $\overline{\varepsilon}\in\mbox{\rm E}_{d-1}(\overline{R})(=\mbox{\rm E}_{d-1}(\overline{\eta}))$ such that $(\overline{\lambda v_{2}},\ldots,\overline{\lambda v_{d}})\overline{\varepsilon}=(\overline{w_{2}},\ldots,\overline{w_{d}}^{(d-1)!}).$ Let $\varepsilon\in\mbox{\rm E}_{d-1}(\eta)\subset\mbox{\rm E}_{d-1}(I)$ be a lift of $\overline{\varepsilon}.$ Therefore $({\lambda v_{2}},\ldots,{\lambda v_{d}}){\varepsilon}=({w_{2}},\ldots,{w_{d}}^{(d-1)!})~\mbox{mod}~(Iv_{1}^{\prime}).$ Thus, we have
$(v_{1},\ldots,v_{d})$ $\underset{\mbox{\rm E}_{d}(R,I)}{\sim}(v_{1}^{\prime},\lambda v_{2},\ldots,\lambda v_{d})$
$\underset{\mbox{\rm E}_{d}(R,I)}{\sim}(w_{1},{w_{2}},\ldots,{w_{d}}^{(d-1)!}),~~~~~~\mbox{where}~w_{1}=v_{1}^{\prime}.$
[…]
Now we consider the problem of injective stability and prove that stably elementary matrices of size $d$ are isotopic to the identity. To prove this we need the following proposition.
Proposition 3.4.
Let $R$ be an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$. Let $v\in\mbox{\rm Um}_{d}(R)$ be such that $v\equiv e_{1}~\mbox{mod}~(s)$ for some $s\in R.$ Then there exists $\sigma\in\mbox{\rm SL}_{d}(R)$ with $\sigma\equiv\mbox{I}_{d}~\mbox{mod}~(s)$ such that $v=e_{1}\sigma.$
$\bf Proof:$}} Let $S=R[X]/(X^{2}-sX).$ Then $S$ is an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$. Since $v\equiv e_{1}~\mbox{mod}~(s),$ let $v=e_{1}+sw$ for some $w\in R^{d}.$ Let $u(X)=e_{1}+Xw.$
Claim : $u(X)\in\mbox{\rm Um}_{d}(S).$
Proof of the claim : Since $v=e_{1}+sw,$ there exist $w_{1}^{\prime},\ldots,w_{d}^{\prime}\in R$ such that
$(1+sw_{1})w_{1}^{\prime}+sw_{2}w_{2}^{\prime}+\cdots+sw_{d}w_{d}^{\prime}=1.$
Suppose to the contrary that $u(X)\notin\mbox{\rm Um}_{d}(S).$ Thus $\langle 1+Xw_{1},Xw_{2},\ldots,Xw_{d}\rangle\subset\mathfrak{p}$ for some $\mathfrak{p}\in\mbox{Spec}(S).$ Therefore
$(1+Xw_{1})w_{1}^{\prime}+\cdots+Xw_{d}w_{d}^{\prime}$ $=w_{1}^{\prime}+X(w_{1}w_{1}^{\prime}+\cdots w_{d}w_{d}^{\prime})$
$=1+(X-s)(w_{1}w_{1}^{\prime}+\cdots w_{d}w_{d}^{\prime})\in\mathfrak{p}.$
Since X(Xs)=0X(X-s)=0 in $S$, $X\in\mathfrak{p}.$ As $1+Xw_{1}\in\mathfrak{p},$ $1\in\mathfrak{p}$ which is not possible. Thus $u(X)\in\mbox{\rm Um}_{d}(S).$
[…]
Theorem 3.5.
Let $R$ be an affine algebra of dimension $d\geq 4$ over $\overline{\mathbb{F}}_{p}$. Let $\alpha\in\mbox{\rm SL}_{d}(R)\cap\mbox{\rm E}_{d+1}(R).$ Then $\alpha$ is isotopic to the identity matrix.
$\bf Proof:$}} Since $\alpha\in\mbox{\rm SL}_{d}(R)\cap\mbox{\rm E}_{d+1}(R)$, there exists a $\beta(T)\in\mbox{\rm SL}_{d+1}(R[T])$ such that
$\beta(0)=I_{d+1}~\mbox{and}~\beta(1)=1\perp\alpha.$
Let $A=R[T],s=T^{2}-T$ and $v=e_{1}\beta(T).$ By Proposition 3.4, there exists $\gamma(T)\in\mbox{\rm SL}_{d+1}(R[T],(s))$ such that $v=e_{1}\gamma(T).$ Therefore $e_{1}\beta(T)\gamma(T)^{-1}=e_{1}.$ Thus
$\beta(T)\gamma(T)^{-1}=(1\perp\eta(T))\Pi_{i=2}^{d+2}E_{i1}(r_{i}),$
where $\eta(T)\in\mbox{\rm SL}_{d+1}(R[T])$ and $r_{i}\in R[T].$ Now,
$\beta(0)\gamma(0)^{-1}=I_{d+1}=1\perp\eta(0),$
and
$\beta(1)\gamma(1)^{-1}=(1\perp\alpha)I_{d+1}=1\perp\eta(1).$
[…]
4. Completion to a symplectic matrix
In this section, we prove that symplectic matrices of size $d$ acts transitively on unimodular rows of length $d$ when $d\equiv 2~(\mbox{mod}~4).$
Lemma 4.1.
Let $R$ be an affine algebra of dimension $d$ over $\overline{\mathbb{F}}_{p}$ and $d\equiv 2~(\mbox{mod}~4).$ Let $v\in\mbox{\rm Um}_{d}(R).$ Assume that $\mbox{\rm SL}_{d+1}(R)\cap\mbox{\rm E}(R)=\mbox{\rm E}_{d+1}(R).$ Then $v$ is completable to a symplectic matrix.
$\bf Proof:$}} Let $d=2$ and $v=(v_{1},v_{2})\in\mbox{\rm Um}_{2}(R).$ Let $w_{1},w_{2}\in R$ be such that $v_{1}w_{1}+v_{2}w_{2}=1.$ Consider
$\alpha=\begin{bmatrix}v_{1}&v_{2}\\ -w_{2}&w_{1}\\ \end{bmatrix}\in\mbox{\rm SL}_{2}(R)=\mbox{\rm Sp}_{2}(R).$
Thus $v$ is completable to a symplectic matrix.
Thus we may assume that $d\geq 6.$ In view of Theorem 3.1, there exists $\varepsilon\in\mbox{\rm E}_{d}(R)$ such that $v\varepsilon=\chi_{d!}(v_{1})$ for some $v_{1}\in\mbox{\rm Um}_{d}(R).$ By Remark 2.5, there exists $\varepsilon_{1}\in\mbox{\rm ESp}_{d}(R)$ such that $v\varepsilon_{1}=\chi_{d!}(v_{1}).$ Let $w_{1}\in\mbox{\rm Um}_{d}(R)$ be such that $\langle v_{1},w_{1}\rangle=1.$ By Lemma 2.6, we have $\varepsilon_{J_{d-1}}\in\mbox{\rm ESp}_{J_{d-1}}(R)$ such that
$S_{d-1}(v,w)\varepsilon_{J_{d-1}}=(I_{2^{d-1}-d}\perp\gamma)\varepsilon_{2}$
for some $\gamma\in\mbox{\rm Sp}_{d}(R)$ and $\varepsilon_{2}\in\mbox{\rm E}_{2^{d-1}}(R).$ Therefore $S_{d-1}(v,w)$ and $\gamma$ are stably elementary equivalent. One knows that $S_{d-1}(v_{1},w_{1})$ and $\beta_{d-1}(v_{1},w_{1})$ are stably elementary equivalent, we have $\beta_{d-1}(v_{1},w_{1})\gamma^{-1}\in\mbox{\rm SL}_{d}(R)\cap\mbox{\rm E}_{d+1}(R).$ By [4, Corollary 3.9], first row of an $1$-stably elementary matrix is elementarily completable. Thus $e_{1}\beta_{d-1}(v_{1},w_{1})\gamma^{-1}=e_{1}\delta^{\prime}$ for some $\delta^{\prime}\in\mbox{\rm E}_{d}(R).$ Now by Remark 2.5, $e_{1}\delta^{\prime}=e_{1}\delta$ for some $\delta\in\mbox{\rm ESp}_{d}(R).$ Therefore
$e_{1}\beta_{d-1}(v_{1},w_{1})\gamma^{-1}=e_{1}\delta.$
[…]
Proposition 4.2.
Let $R$ be an affine algebra of dimension $d$ over $\overline{\mathbb{F}}_{p}$ and $d\equiv 2~\mbox{mod}~4.$ Let $v\in\mbox{\rm Um}_{d}(R)$ be such that $v\equiv e_{1}~\mbox{mod}~(s)$ for some $s\in R.$ Assume that $\mbox{\rm SL}_{d+1}(R)\cap\mbox{\rm E}(R)=\mbox{\rm E}_{d+1}(R).$ Then there exists $\sigma\in\mbox{\rm Sp}_{d}(R)$ be such that $\sigma\equiv\mbox{I}_{d}~(\mbox{mod}~s)$ and $v=e_{1}\sigma.$
[…]
The next result shows that a stably elementary symplectic matrix over $R$ is isotopic to the identity matrix under certain conditions when $R$ is an affine algebra of dimension $d$ over $\overline{\mathbb{F}}_{p}.$
Theorem 4.3.
Let $R$ be an affine algebra of dimension $d$ over $\overline{\mathbb{F}}_{p}.$ Assume that $1/(d+1)!\in\overline{\mathbb{F}}_{p}$ and $d\equiv 1~(\mbox{mod}~4).$ Further assume that $\mbox{\rm SL}_{d+1}(R)\cap\mbox{\rm E}(R)=\mbox{\rm E}_{d+1}(R).$ Let $\sigma\in\mbox{\rm Sp}_{d-1}(R)\cap\mbox{\rm ESp}_{d+1}(R).$ Then $\sigma$ is isotopic to the identity.
$\bf Proof:$}} Since $\sigma\in\mbox{\rm Sp}_{d-1}(R)\cap\mbox{\rm ESp}_{d+1}(R)$, there exists $\delta(T)\in\mbox{\rm ESp}_{d+1}(R[T])$ be such that $\delta(1)=I_{2}\perp\sigma,\delta(0)=\mbox{I}_{d+1}.$ Let $v(T)=e_{1}\delta(T)\in\mbox{\rm Um}_{d+1}(R[T],(T^{2}-T)).$ By Proposition 4.2, $v(T)=e_{1}\alpha(T)$ for some $\alpha(T)\in\mbox{\rm Sp}_{d+1}(R[T])$ and $\alpha(T)\equiv\mbox{I}_{d+1}~\mbox{mod}~(T^{2}-T).$ Thus $e_{1}\delta(T)=e_{1}\alpha(T).$ Therefore one has $e_{1}\delta(T)\alpha(T)^{-1}=e_{1}.$ Now note that $\delta(T)\alpha(T)^{-1}$ is an isotopy of $\mbox{I}_{2}\perp\sigma$ to $\mbox{I}_{d+1}.$ The matrix $\delta(T)\alpha(T)^{-1}$ looks like
$\delta(T)\alpha(T)^{-1}=\begin{bmatrix}1&0&0\\ \ast&1&\ast\\ \ast&0&\eta(T)\\ \end{bmatrix}$
[…]